[ -x + 3y = -9 ]
[ 8x-4y = 32 ]

Answer by TimothyLamb(4379)

-x + 3y = -9
8x - 4y = 32
---
-(8)x + (8)3y = (8)-9
8x - 4y = 32
---
-8x + 24y = -72
8x - 4y = 32
---
add:
20y = -40
---
y = -2
x = 3
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php